OpenStudy (anonymous):

plz help asap What is the volume of the prism? The triangular base has a base of 9 yards and 6 yards the height of the prism is 16 yards

OpenStudy (anonymous):

answer choices 228 cubic yards 432 cubic yards 576 cubic yards 864 cubic yards

ganeshie8 (ganeshie8):

volume of prism = (area of base) x height

ganeshie8 (ganeshie8):

height = 16 area of base = ?

ganeshie8 (ganeshie8):

lol

OpenStudy (anonymous):

um.. 15?

ganeshie8 (ganeshie8):

no, u need to use the area of triangle formula

ganeshie8 (ganeshie8):

area of base = 1/2 bh = 1/2 9 x 6 = 27

ganeshie8 (ganeshie8):

multiply this base area with the height, you wil get volume

ganeshie8 (ganeshie8):

:) i wrote the formula above, and solved the area of base for you

OpenStudy (anonymous):

come back rex the dinosaur i need you

ganeshie8 (ganeshie8):

im here

OpenStudy (anonymous):

okay so all I do is multiple the area and base with the height

ganeshie8 (ganeshie8):

all you do is multiply the area of* base wid height of prism

OpenStudy (anonymous):

so 16x27

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

so thats 432

ganeshie8 (ganeshie8):

looks correct :)

OpenStudy (anonymous):

yes my awesomeness has struck again

ganeshie8 (ganeshie8):

you're fully awesome today ;)

OpenStudy (anonymous):

i know thats right boom boom fire power!